Hocking Corporations comparative balance sheet appears below: Hocking Corporation Comparative Balance Sheet Ending Balance Beginning Balance Assets Current assets Cash and cash equivalents Accounts recelvable Inventory Prepald expenses Total current assets Property, plant, and equipment Less accumulated depreclation Net property, plant, and equipment Total assets Liabilities and stockholders equity Current liablities Accounts payable Accrued llabllitles Income taxes payable Total current llabllitles Bonds payable Total llabilitles Stockholders equity Common stock Retalned earnings Total stockholders equity Total liablities and stockholders equity $35,000 20,000 66,000 12,800 133,800 422,000 211,000 211,000 $27,300 24,000 62,000 12 8010400 123.700 402,000 180,900 221,100 $13,000 63,000 79,000 155,000 81,000 20 236,000 $19,400 56,000 77,000 152,400 88,000 240,40 36,000 68,400 104,400 $344,800 30,000 78,800 108,800 $344,800 The companys net Income (loss) for the year was $(5,200) and its cash dividends were $5,200. It did not sell or retire any property, plant, and equlpment during the year. The company uses the Indirect method to determine the net cash provided by operating activities. The companys net cash provided by operating activities Is O $19,900 O $30,300 O $34,300 O $19,700
Answer #1
Cash Flows from Operating Activities:
Net income -5200
Adjustments to reconcile net income to
Net cash provided by operating activities
Depreciation expense 30100 =211000-180900
Decrease in Accounts receivable 4000 =24000-20000
Increase in inventory -4000 =62000-66000
Increase in prepaid expenses -2400 =10400-12800
Increase in accounts payable 6400 =19400-13000
Decrease in accrued liabilities -7000 =56000-63000
Decrease in income taxes payable -2000 =77000-79000
25100
Net cash provided by operating activities 19900
Option 1 $19,900 is correct